# Improving Algorithmic Efficiency using Cryptography

### v1

@@ 3 Trapdoored Matrices from LPN
since one first cuts $\mathbf{v}$ into $n/k$ pieces of length $k$ each, computes $n/k$ matrix-vector multiplications in $k$ dimensions followed by $n/k$ additions of $k$-dimensional vectors to compute $\mathbf{B}\mathbf{v}$, and another $n/k$ matrix-vector multiplications in $k$ dimensions to compute $\mathbf{A}\mathbf{B}\mathbf{v}$. Computing $\mathbf{E}\mathbf{v}$ takes another $pn^{2}$ operations.
@@
    $\mathbf{B}_{i}:=[\mathbf{B}_{i,1}||\mathbf{B}_{i,2}||\ldots||\mathbf{B}_{i,n_{i}/n_{i+1}}]\in\mathbb{F}^{n_{i+1}\times n_{i}}~{};$
@@
$T(n_{i})$ $=\frac{2n_{i}}{n_{i+1}}\cdot T(n_{i+1})+O(n_{i}+pn_{i}^{2})$
$=2n_{i}^{\epsilon}\cdot T(n_{i}^{1-\epsilon})+O\big{(}n_{i}^{2-\delta(1-\epsilon)}\big{)}$
@@
where the first inequality is by approximating the geometric sum $\sum_{j=0}^{\ell}(1-\epsilon)^{j}$ by $1/\epsilon$. Under the subexponential LPN assumption, we can only go as far as getting $n_{\ell}=n^{(1-\epsilon)^{\ell}}$ to be $\mathsf{poly}(\log n)$, which means this sum will be $n\cdot\mathsf{poly}(\log n)$. The second term of the sum becomes
@@
where $c:=2-\delta\cdot(1-\epsilon)>1$. Note the exponents decrease monotonically since $c<1$, and so the sum is dominated by the first term. Setting $\delta\approx 1$ and $\epsilon\approx 0$ gives us $c\approx 1$, and so the sum is $O(n\cdot\mathsf{poly}(\log n))$.
@@
If we only assume the polynomial LPN assumption, on the other hand, we will stop the recursion at $n_{\ell}=n^{\gamma}$ for an arbitrarily small constant $\gamma>0$. Thus will result in a complexity $T(n)$ that is $n^{1+\gamma}$ for an arbitraily small positive constant $\gamma$.
@@
4 Trapdoored Matrices from McEliece
@@
$O(\bigg{(}\frac{n}{k}\cdot(k^{2}+n)+\frac{n}{k}\cdot n\bigg{)}=O\bigg{(}nk+\frac{n^{2}}{k}\bigg{)}$
@@ 5 Trapdoored Matrices over the Reals
We introduce a conjecture on Kac’s random walks, yielding a simple construction of trapdoored matrices over the reals.
@@
We define a random walk starting from $Q_{0}=I_{n}$ in which every step consists of a multiplication with a matrix $R_{i,j,\theta}$ where the indices $i,j$ and the angle $\theta$ are drawn uniformly. That is $Q_{t}\leftarrow R_{i_{t},j_{t},\theta_{t}}$. The resulting matrix is the state $Q_{T}$ of the random walk after $T$ steps. Essentially, in each step we perform a random rotation in low-dimension. It is thus reasonable to expect that after enough steps we will converge to a random “rotation” of the entire space.
@@
Understanding the mixing time of the Kac’s random walk over the sphere (that is, considering the evolution of only a single column of $Q_{t}$, or equivalently of $Q_{t}\cdot e_{1}$) already required significant mathematical effort. In a sequence of breakthrough papers [Jan01, MM13, PS17] it was proven that when $T=\Theta(n\log n)$ the distribution of the resulting state $Q_{t}\cdot e_{1}$ is already approaching the uniform distribution of $S^{n-1}$, even in the total variation distance. This is the best possible mixing time due to a simple lower coupon collector-style lower bound.
@@
Understanding the mixing time of the entire transformation $Q_{t}$ is even more challenging. Although it is easy to show that the chain converges to the uniform distribution in the special orthogonal group $SO(n)$, only in 2012, Jiang [Jia12] showed that the chain mixes in polynomial time. The bounds have since been improved and it is currently known that the mixing time is between $\Omega(n^{2})\leq T\leq O(n^{4}\log n)$.
@@
We pose the following natural conjecture, stating that after $n\;\mathsf{poly}\log n$ the resulting state $Q_{T}$ should already be computationally indistinguishable from the uniform distribution over $SO(n)$. As stated above, after that number of steps any small set of columns is distributed as i.i.d. samples from $S^{n-1}$, yet the distribution of the entire matrix is still not statistically close to uniform. We are not aware of previous conjectures or works considering the computational psuedo-randomness of these walks.
Conjecture.
@@
There exists some $T=n\;\mathsf{poly}\log(n)$ such that the distribution of $Q_{T}$ is computationally indistinguishable from the distribution $U\left(SO\left(n\right)\right)$.
@@
We note that many existing hardness assumptions can also be phrased as running a mixing Markov chain for fewer steps than necessary for statistical convergence and conjecturing that the resulting state is pseudo-random. Further supporting this conjecture is a recent result [JPS+22] in which it was shown that the transformation $Q_{T}$ with $T=\Theta(n\log n)$ suffices for recovering the JL Lemma. This is an intricate unconditional proof that $Q_{T}$ preserves all pair-wise distances of a point-set w.h.p., recovering the previous results on Fast JL Transforms [AL13]. We conjecture that the same distribution is in fact psuedorandom, and would thus suffice for the far more general applications as well.

### v2

@@ 3.2 Construction from LPN
since one first cuts $\mathbf{v}$ into $n/k$ pieces of length $k$ each, computes $n/k$ matrix-vector multiplications in $k$ dimensions followed by $n/k$ additions of $k$-dimensional vectors to compute $\mathbf{B}\mathbf{v}$, and another $n/k$ matrix-vector multiplications in $k$ dimensions to compute $\mathbf{A}\mathbf{B}\mathbf{v}$. Computing $\mathbf{E}\mathbf{v}$ takes another $pn^{2}$ operations, and adding the two takes $n$ operations.
@@
    where $[\mathbf{A}\sslash\mathbf{B}]$ refers to concatenating the matrices $\mathbf{A}$ and $\mathbf{B}$ vertically.
@@
    $\mathbf{B}_{i}:=[\mathbf{B}_{i,1}\|\mathbf{B}_{i,2}\|\ldots\|\mathbf{B}_{i,n_{i}/n_{i+1}}]\in\mathbb{F}^{n_{i+1}\times n_{i}}~{};$
    where where $[\mathbf{A}\|\mathbf{B}]$ refers to concatenating the matrices $\mathbf{A}$ and $\mathbf{B}$ horizontally.
@@
$T(n_{i})$ $=\frac{2n_{i}}{n_{i+1}}\cdot T(n_{i+1})+O(n_{i}+p_{i}n_{i}^{2})$
$=2n_{i}^{\epsilon}\cdot T(n_{i}^{1-\epsilon})+O\big{(}n_{i}^{\mathsf{max}(1,2-\delta(1-\epsilon))}\big{)}$
@@
where the first inequality is by upper-bounding the geometric sum $\sum_{j=0}^{\ell}(1-\epsilon)^{j}$ by $1/\epsilon$. Under the subexponential LPN assumption, we can only go as far as setting $n_{\ell}=n^{(1-\epsilon)^{\ell}}$ to be $\mathsf{poly}(\log n)$, which means this sum will be $n\cdot\mathsf{poly}(\log n)$. The second term of the sum becomes
@@
where $c:=\mathsf{max}(1,2-\delta\cdot(1-\epsilon))>1$. Note the exponents decrease monotonically since $c>1$, and so the sum is dominated by the first term. Setting $\delta\approx 1$ and $\epsilon\approx 0$ gives us $c\approx 1$, and so the sum is $O(n\cdot\mathsf{poly}(\log n))$.
@@
If we only assume the polynomial LPN assumption, on the other hand, we will stop the recursion at $n_{\ell}=n^{\gamma}$ for an arbitrarily small constant $\gamma>0$. Thus will result in a complexity $T(n)$ that is $n^{1+O(\gamma)}$ for an arbitraily small positive constant $\gamma$.
@@
Left and Right Multiplication.
It is easy to see that our pseudorandom matrix $\mathbf{M}=\mathbf{A}\mathbf{B}+\mathbf{E}$ supports both left and right multiplication in near linear time as long as the constituent submatrices of $\mathbf{A}$ and $\mathbf{B}$ have the same property recursively. It is unclear, though, whether one can modify the construction to support quick multplication of a vector by $\mathbf{M}^{-1}$ as well as $\mathbf{M}$.
4 Trapdoored Matrices from the McEliece Assumption
4.1 The McEliece Assumption Family
where $\mathsf{Perm}(n)$ is the set of all $n\times n$ permutation matrices.111 For a reader familiar with the McEliece encryption scheme, we remark that the assumption as formulated above is sufficent for the security of the McEliece cryptosystem (in conjunction with the LPN assumption over $\mathbb{F}$), but is not necessary for it.
McEliece formulated the assumption with $\mathcal{C}$ being the Goppa code [McE78]; this still remains unbroken and is the basis of candidate encryption schemes submitted to the NIST post-quantum cryptography standardization process [Cla23]. The McEliece assumption w.r.t. “quasi-cyclic” codes has also been considered in the literature [MTSB13], and is the basis of a different NIST encryption candidate [bik17]. All these choices of codes permit near-linear time encoding: that is, given a vector $\mathbf{v}\in\mathbb{F}^{k}$, one can compute $\mathbf{G}\mathbf{v}\in\mathbb{F}^{n}$ in time $\tilde{O}(n+k)$, as opposed to the trivial $O(nk)$. On the other hand, when $\mathcal{C}$ is either the Reed-Solomon or the Reed-Muller code, the assumption is broken via an attack of Sidelnikov and Shestakov [SS92]. In the case of Goppa codes, despite recent attempts at cryptanalysis (see, e.g. [FGO+13, COT14, CMT23, BMT24]), the best known attack from earlier this year incurs very slightly subexponential complexity [Ran24, Theorem 3].
Our assumption will be that the McEliece assumption holds w.r.t. a code family that permits near-linear-time encoding, e.g. the Goppa or quasi-cyclic codes [Cla23, bik17, MTSB13]. As in the case of the LPN assumption, we can assume either the (conservative) polynomial or the (aggressive) subexponential versions of the assumption, with different results.
4.2 Construction from McEliece
@@
$O\bigg{(}\frac{n}{k}\cdot(k^{2}+n)+\frac{n}{k}\cdot n\bigg{)}=O\bigg{(}nk+\frac{n^{2}}{k}\bigg{)}$
@@ 5 Trapdoored Matrices over the Reals
5.1 Construction from Kac’s Random Walks
We introduce a conjecture on Kac’s random walks, which immediately yields a simple construction of trapdoored matrices over the reals for the Haar-measure (i.e., the uniform distribution over rotation matrices). We then use this family of trapdoored matrices to construct another family for any Haar-invariant distribution.
@@ 5.1 Construction from Kac’s Random Walks
We define a random walk starting from $Q_{0}=I_{n}$ in which every step consists of a multiplication with a matrix $R_{i,j,\theta}$ where the indices $i,j$ and the angle $\theta$ are drawn uniformly. That is $Q_{t}\leftarrow R_{i_{t},j_{t},\theta_{t}}$. The resulting matrix is the state $Q_{T}$ of the random walk after $T$ steps. Essentially, in each step, we perform a random rotation in low-dimension. It is thus reasonable to expect that after enough steps we will converge to a random “rotation” of the entire space.
@@
Understanding the mixing time of the Kac’s random walk over the sphere (that is, considering the evolution of only a single column of $Q_{t}$, or equivalently of $Q_{t}\cdot e_{1}$) already required significant mathematical effort. In a sequence of breakthrough papers [Jan01, MM13, PS17] it was proven that when $T=\Theta(n\log n)$ the distribution of the resulting state $Q_{t}\cdot e_{1}$ approaches the uniform distribution of $S^{n-1}$, even in the total variation distance. This is the best possible mixing time due to a simple lower coupon collector-style lower bound.
@@
Understanding the mixing time of the entire transformation $Q_{t}$ is even more challenging. Although it is easy to show that the chain converges to the uniform distribution in the special orthogonal group $SO(n)$, only in 2012, Jiang [Jia12] showed that the chain mixes in polynomial time. The bounds have since been improved and it is currently known that the mixing time $T(n)$ is between $\Omega(n^{2})\leq T(n)\leq O(n^{4}\log n)$.
@@
We pose the following natural conjecture, stating that after $n\cdot\mathsf{poly}(\log n)$ steps, the resulting state $Q_{T}$ should already be computationally indistinguishable from the uniform distribution over $SO(n)$. As stated above, after that number of steps any small set of columns is distributed as i.i.d. samples from $S^{n-1}$, yet the distribution of the entire matrix is still not statistically close to uniform. We are not aware of previous conjectures or works considering the computational pseudo-randomness of these walks.
Conjecture 5.1.
@@
There exists some $T=n\cdot\mathsf{poly}\log(n)$ such that the distribution of $Q_{T}$ is computationally indistinguishable from the Haar-measure over $SO\left(n\right)$.
@@
We note that many existing hardness assumptions can also be phrased as running a mixing Markov chain for fewer steps than necessary for statistical convergence and conjecturing that the resulting state is pseudo-random. Further supporting this conjecture is a recent result [JPS+22] in which it was shown that the transformation $Q_{T}$ with $T=\Theta(n\log n)$ suffices for recovering the JL Lemma. This is an intricate unconditional proof that $Q_{T}$ preserves all pair-wise distances of a point-set w.h.p., recovering the previous results on Fast JL Transforms [AL13]. We conjecture that the same distribution is in fact pseudorandom, and would thus suffice for the far more general applications as well.